OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1 medal awarded

OpenStudy (anonymous):

3x-4y=-24 x+y = -1 3x-4y=-24 - 3x +3y = -3 ------------- -7y = -21 7 = 3

OpenStudy (anonymous):

I mean y = 3

OpenStudy (anonymous):

See if you can solve from there :)

OpenStudy (anonymous):

Just substitute y=3 back into the 2nd equation.

OpenStudy (anonymous):

-4

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

The local zoo is filling two water tanks for the elephant exhibit. One water tank contains 34 gal of water and is filled at a constant rate of 4 gal/h. The second water tank contains 30 gal of water and is filled at a constant rate of 2 gal/h. Explain. Let x = the number of hours the tanks are filling and let y = the number of gallons in the tank. What are the constraints on the system? do u know this

OpenStudy (anonymous):

I'm sorry, no. But You could post this up and see if anyone knows?

OpenStudy (anonymous):

Which ordered pair is a solution of the inequality? y>=4x - 5

OpenStudy (anonymous):

\[y \ge4x-5\]?

OpenStudy (anonymous):

What are the options?

OpenStudy (anonymous):

can u solve it without or no because i wanted to see if any answers matched yours

OpenStudy (anonymous):

You can't solve without for this question

OpenStudy (anonymous):

A. (3, 4) B. (2, 1) C. (3, 0) D. (1, 1)

OpenStudy (anonymous):

i agree

OpenStudy (anonymous):

You just plug in these values into the equation and see if it makes sense.

OpenStudy (anonymous):

A

OpenStudy (anonymous):

e.g \[4 \ge 4(3)-5 ?\]

OpenStudy (anonymous):

That isn't correct.

OpenStudy (anonymous):

the third \[0 \ge 3(3) -5?\]

OpenStudy (anonymous):

I think it's D

OpenStudy (anonymous):

the fourth \[1 \ge 4(1)-5?\]

OpenStudy (anonymous):

D is correct

OpenStudy (anonymous):

a and c is wrong

OpenStudy (anonymous):

Sorry, the second is 1> 4(2) -4

OpenStudy (anonymous):

-5*

OpenStudy (anonymous):

which isn't correct. So, only D can be the answer.

OpenStudy (anonymous):

Graph the inequality. y < 3x - 5

OpenStudy (anonymous):

Plug in values for y and x, eg. when y = 0, x =? and when x= 0, y =? Then you have two points and can graph it. Or, get more values and graph it too

OpenStudy (anonymous):